# A Training-Free Framework for Precise Mobile Manipulation of Small Everyday Objects

### v1

@@ V Experiments
Our experiments are designed to test a) the extent to which open loop execution is an issue for precise mobile manipulation tasks, b) how effective are blind proprioceptive correction techniques, c) do object detectors and point trackers perform reliably enough in wrist camera images for reliable control, d) is occlusion by the end-effector an issue and how effectively can it be mitigated through the use of video in-painting models, and e) how does our proposed SVM methodology compare to large-scale imitation learning?
@@ V-A Tasks and Experimental Setup
We work with the Stretch RE2 robot. Stretch RE2 is a commodity mobile manipulator with a 5DOF arm mounted on top of a non-holomonic base. We upgrade the robot to use the Dex Wrist 3, which has an eye-in-hand RGB-D camera (Intel D405). We consider 3 task families for a total of 6 different tasks: a) holding a knob to pull open a cabinet or drawer, b) holding a handle to pull open a cabinet, and c) pushing on objects (light buttons, books in a book shelf, and light switches). Our focus is on generalization. Therefore, we exclusively test on previously unseen instances, not used during development in any way. Figure 5 shows the instances that we test on.
All tasks involve some precise manipulation, followed by execution of a motion primitive. For the pushing tasks, the precise motion is to get the end-effector exactly at the indicated point and the motion primitive is to push in the direction perpendicular to the surface and retract the end-effector upon contact. The robot is positioned such that the target position is within the field of view of the wrist camera. A user selects the point of pushing via a mouse click on the wrist camera image. The goal is to push at the indicated location. Success is determined by whether the push results in the desired outcome (light turns on / off or book gets pushed in). The original rubber gripper bends upon contact, we use a rigid known tool that sticks out a bit. We take the geometry of the tool into account while servoing.
For the opening articulated object tasks, the precise manipulation is grasping the knob / handle, while the motion primitive is the whole-body motion that opens the cupboard. Computing and executing this full body motion is difficult. We adopt the modular approach to opening articulated objects (MOSART) from Gupta et al. [47] and invoke it after the gripper has been placed around the knob / handle. The whole tasks starts out with the robot about 1.5m way from the target object, with the target object in view from robot’s head mounted camera. We use MOSART to compute articulation parameters and convey the robot to a pre-grasp location with the target handle in view of the wrist camera. At this point, SVM (or baseline) is used to center the gripper around the knob / handle, before resuming MOSART: extending the gripper till contact, close the gripper, and play rest of the predicted motion plan. Success is determined by whether the cabinet opens by more than $60^{\circ}$ or the drawer is pulled out by more than $24cm$, similar to the criteria used in [47].
@@
Table I: Execution Success Rates. We compare Servoing with Vision Models (SVM) to a previous system (MOSART [47]), open loop execution using target computed in the wrist camera, and a version of SVM without inpainting. Tasks require precise control and open loop execution fails. MOSART’s contact correction works for handles but struggles with knobs and it can’t handle user-clicked targets. In-painting matters for semantic targets.
Semantic Targets User-clicked Targets Total
Knobs (Cabinets) Knobs (Drawers) Handles (Cabinets) Light Button Book Light Switch
MOSART [47] 1/7 0/7 6/7 - - - 7/28
Open Loop (Eye-in-Hand) 2/7 0/7 6/7 0/3 2/2 0/2 10/28
Ours w/o inpainting 6/7 3/7 4/7 3/3 2/2 2/2 20/28
Ours 6/7 5/7 6/7 3/3 2/2 2/2 24/28
Figure 6: Comparison of SVM with the open loop (eye-in-hand) baseline for opening a cabinet with a knob. Slight errors in getting to the target cause the end-effector to slip off, leading to failure for the baseline, where as our method is able to successfully complete the task.
Knobs Handle Total
Cabinets Drawer Cabinets
RUM [7] 0/3 1/4 1/3 2/10
SVM (Ours) 2/3 2/4 3/3 7/10
Table II: Comparison of SVM vs. RUM [7], a recent large-scale end-to-end imitation learning method trained on 1200 demos for opening cabinets and 525 demos for opening drawers across 40 different environments. Our evaluation spans objects from three environments across two buildings.
@@ V-B Baselines
We compare against three other methods for the precise manipulation part of these tasks.
V-B1 Open Loop (Eye-in-Hand)
To assess the precision requirements of the tasks and to set it in context with the manipulation capabilities of the robot platform, this baseline uses open loop execution starting from estimates for the 3D target position from the first wrist camera image.
V-B2 MOSART [47]
The recent modular system for opening cabinets and drawers [47] reports impressive performance with open-loop control (using the head camera from 1.5m away), combined with proprioception-based feedback to compensate for errors in perception and control when interacting with handles. We test if such correction is also sufficient for interacting with knobs. Note that such correction is not possible for the smaller buttons and pliable books.
V-B3 SVM (no inpainting)
To understand how much of an issue occlusion due to the end-effector is during manipulation, we ablate the use of inpainting.
V-B4 Robot Utility Models (RUM) [7]
For the opening articulated object tasks, we also compare to Robot Utility Models (RUM), a closed-loop imitation learning method recently proposed by Etukuru et al. [7]. RUM is trained on a substantial dataset comprising expert demonstrations, including 1,200 instances of cabinet opening and 525 of drawer opening, gathered from roughly 40 different environments. This dataset stands as the most extensive imitation learning dataset for articulated object manipulation to date, establishing RUM as a strong baseline for our evaluation.
Similar to our method, we use MOSART to compute articulation parameters and convey the robot to a pre-grasp location with the target handle in view of the wrist camera. One of the assumptions of RUM is a good view of the handle. To benefit RUM, we try out three different heights of the wrist camera, and report the best result for RUM.
Figure 7: SVM vs. open loop (eye-in-hand) baseline for pushing on user-clicked points. Slight errors in getting to the target cause failure, where as SVM successfully turns the lights off. Note the quality of CoTracker’s track ( blue dot).
Figure 8: Comparison of SVM with and without inpainting. Erroneous detection without inpainting causes execution to fail, where as with inpainting the target is correctly detected leading to a successful grasp and a successful execution.
@@ V-C Results
Table I presents results from our experiments. Our training-free approach SVM successfully solves over 85% of task instances that we test on. As noted, all these tests were conducted on unseen object instances in unseen environments that were not used for development in any way. We discuss our key experimental findings below.
V-C1 Closing the loop is necessary for these precise tasks
While the proprioception-based strategies proposed in MOSART [47] work out for handles, they are inadequate for targets like knobs and just don’t work for tasks like pushing buttons. Using estimates from the wrist camera is better, but open loop execution still fails for knobs and pushing buttons.
V-C2 Vision models work reasonably well even on wrist camera images
@@
Inpainting works well on wrist camera images (see Figure 2 and Figure 3). Closing the loop using feedback from vision detectors and point trackers on wrist camera images also work well, particularly when we use in-painted images. See some examples detections and point tracks in Figure 6 and Figure 7. Detic [2] was able to reliably detect the knobs and handles and CoTracker [3] was able to successfully track the point of interaction letting us solve 24/28 task instances.
V-C3 Erroneous detections without inpainting hamper performance on handles and our end-effector out-painting strategy effectively mitigates it
@@
As shown in Figure 8, presence of the end-effector caused the object detector to miss fire leading to failed execution. Our out painting approach mitigates this issue leading to a higher success rate than the approach without out-painting. Interestingly, CoTracker [3] is quite robust to occlusion (possibly because it tracks multiple points) and doesn’t benefit from in-painting.
V-C4 Closed-loop imitation learning struggles on novel objects
@@
As presented in Table II, SVM significantly outperforms RUM in a paired evaluation on unseen objects across three novel environments. A common failure mode of RUM is its inability to grasp the object’s handle, even when it approaches it closely. Another failure mode we observe is RUM misidentifying keyholes or cabinet edges as handles, also resulting in failed grasp attempts. These result demonstrate that a modular approach that leverages the broad generalization capabilities of vision foundation models is able to generalize much better than an end-to-end imitation learning approach trained on 1000+ demonstrations, which must learn all aspects of the task from scratch.
@@ VI Discussion
In this paper, we describe SVM, a training-free framework for precise manipulation tasks that involve reaching a precise interaction site followed by execution of a primitive motion. Strong vision models help us mitigate issues caused by occlusion by the end-effector thereby enabling the use of off-the-shelf open-vocabulary object detectors and point trackers to estimate servoing targets during execution. Use of strong off-the-shelf models also provides broad generalization for perception while the use of servoing provides robust control. This enables SVM to solve tasks in novel environments on novel objects, obtaining a 85% zero-shot success rate across 4 precise mobile manipulation tasks. Most surprisingly, even though SVM is modular, it outperforms the strong end-to-end imitation learning system RUM [7] that was trained on 1000+ demonstrations. This is particularly striking as imitation learning is the tool of choice for precise tasks requiring closed-loop execution.

### v2

@@ V Experiments
We conduct experiments in 10 previously unseen environments across 6 previously unseen buildings including 72 different object instances (see Figure 4). Our experiments are designed to test a) the extent to which open-loop execution is an issue for precise mobile manipulation tasks, b) the effectiveness of blind proprioceptive correction techniques, c) whether object detectors and point trackers perform reliably in wrist camera images, d) the impact of end-effector occlusion and the efficacy of video inpainting in mitigating it, and e) how our proposed SVM framework compares to large-scale imitation learning.
@@ V-A Tasks and Experimental Setup
We use the Stretch RE2, a commodity mobile manipulator with a 5DOF arm on a non-holonomic base, upgraded with the Dex Wrist 3 and an eye-in-hand RGB-D camera (Intel D405). We evaluate SVM across 3 task families: a) grasping knobs to open drawers/cabinets, b) grasping handles to open cabinets, and c) pushing objects (buttons, books, switches). Our focus is on generalization. Therefore, we exclusively test on previously unseen instances, not used during development.
All tasks involve precise manipulation, followed by execution of a motion primitive. For the pushing tasks, the precise motion is to get the end-effector exactly at the indicated point and the motion primitive is to push in the direction perpendicular to the surface, before retracting the end-effector upon contact. The robot is positioned such that the target position is within the field of view of the wrist camera. A user selects the point of pushing via a mouse click on the wrist camera image. The goal is to push at the indicated location. Success is determined by whether the push results in the desired outcome (light turns on / off or book gets pushed in). Since the rubber gripper bends on contact, we attach a rigid tool and account for its geometry during servoing.
For the opening articulated object tasks, the precise manipulation involves grasping a knob or handle, followed by whole-body motion to open the cabinet. We adopt MOSART [39] for this motion. Each episode begins with the robot $\approx 1.5m$ from the object, visible from the head camera. MOSART computes articulation parameters and navigates to a pre-grasp pose. SVM (or baseline) then centers the gripper around the knob/handle, after which MOSART resumes: it extends until contact, closes the end-effector, and finally executes the predicted motion plan. Following [39], success is defined as opening the cabinet by over $60^{\circ}$ or the drawer by more than 24cm.
@@ V-B Baselines
We compare against four other methods for the precise manipulation part of these tasks.
Figure 5: SVM vs. Open-Loop (Eye-in-Hand) baseline. (top) In opening a cabinet with a knob, slight errors in getting to the target cause the end-effector to slip off, leading to failure for the baseline, whereas our method is able to successfully complete the task. (bottom) Slight errors in getting to the target cause failure, whereas SVM successfully turns the lights off. Note the high quality of CoTracker’s track ( blue dot).
MOSART [39]. The recent modular system for opening cabinets and drawers [39] reports impressive performance with open-loop control (using the head camera from 1.5m away), combined with proprioception-based feedback to compensate for errors in perception and control when interacting with handles. We test if such correction is also sufficient for interacting with knobs. Note that such correction is not possible for the smaller buttons and pliable books.
Open-Loop (Eye-in-Hand). To assess the precision requirements of the tasks and to set it in context with the manipulation capabilities of the robot platform, this baseline uses open-loop execution starting from estimates for the 3D target position from the first wrist camera image.
SVM (no inpainting). To assess the negative impact of the end-effector’s presence in the wrist camera image during manipulation, we ablate the use of inpainting.
Robot Utility Models (RUM) [7]. For the opening articulated object tasks, we also compare to Robot Utility Models (RUM), a closed-loop imitation learning method recently proposed by Etukuru et al. [7]. RUM is trained on a substantial dataset comprising expert demonstrations, including 1,200 instances of cabinet opening and 525 of drawer opening, gathered from roughly 40 different environments. This dataset stands as the most extensive imitation learning dataset for articulated object manipulation to date, establishing RUM as a strong baseline for our evaluation.
For RUM, similar to SVM, we use MOSART to compute articulation parameters and convey the robot to a pre-grasp location with the target handle in view of the wrist camera. One of RUM’s assumptions is a good view of the handle. To benefit RUM, we try out three different heights of the wrist camera, and report the best result for RUM.
@@ V-C Results
Table I and Table II present results from our experiments. We discuss our key experimental findings from our large-scale evaluation spanning real world experiments across 10 novel environments below.
Vision foundation models with servoing enable generalizable and precise mobile manipulation. Our approach SVM successfully solves 71% of unseen object instances in unseen environments that were not used for development.
@@
Vision foundation models work reasonably well even on wrist camera images. Inpainting works well on wrist camera images (see Figure 2). Closing the loop using feedback from vision detectors and point trackers on wrist camera images also work well, particularly when we use inpainted images. See some example detections and point tracks in Figure 5. Detic [2] was able to reliably detect the knobs and handles, and CoTracker [3] was able to successfully track the point of interaction letting us solve 30/42 task instances.
Closing the loop is necessary for these precise tasks. While the proprioception-based strategies proposed in MOSART [39] work out for handles, they are inadequate for targets like knobs and just do not work for tasks like pushing buttons. Using estimates from the wrist camera is better, but open-loop execution still fails for knobs and pushing buttons (see Figure 5).
@@
Erroneous detections without inpainting hamper performance on handles and our end-effector out-painting strategy effectively mitigates it. As shown in Figure 2, presence of the end-effector caused the object detector to misfire, leading to failed execution. Our out-painting approach mitigates this issue, leading to a higher success rate than the approach without out-painting. Interestingly, CoTracker [3] is quite robust to occlusion and does not benefit from inpainting.
Table II: Comparison of SVM vs. RUM [7], a recent large-scale end-to-end imitation learning method trained on 1200 demos for opening cabinets and 525 demos for opening drawers across 40 different environments. Our evaluation spans objects from 9 previously unseen environments across 6 buildings. SVM significantly outperforms RUM. Note that RUM was not trained on knobs.
Handle Knobs Total
Cabinets Drawer Cabinets Drawer
RUM [7] 2/8 4/7 0/7 1/8 7/30
SVM (Ours) 6/8 7/7 4/7 5/8 22/30
Figure 6: SVM vs RUM [7]. We show roll-outs of both SVM and RUM: at the start of the episode, at the start of the closed-loop grasping, just before closing the gripper, and during manipulation.
Figure 7: SVM Failure Modes. We characterize the failure modes of SVM of all the 20 failures we encountered in Table I and Table II. On the right, we show some examples of these failure modes.
@@
Closed-loop imitation learning struggles on novel objects. As presented in Table II, SVM significantly outperforms RUM in a paired evaluation on unseen objects across 9 novel environments spanning 6 buildings. For handles, RUM succeeds on 6/15 handles vs. SVM succeeds on 13/15 handles. Although RUM was not explicitly trained on knobs, its exposure to a wide variety of handles during training suggested potential for generalization, which motivated our evaluation on knobs as well. While RUM achieves a non-zero success rate on knobs, it struggles relative to SVM (see Figure 6). One failure mode of RUM is failing to grasp the handle despite approaching it closely. These results demonstrate that a modular approach that leverages the broad generalization capabilities of vision foundation models is able to generalize much better than end-to-end imitation learning trained on 1000+ demonstrations, which must learn all aspects of the task from scratch.
SVM Failure Modes. Across 20 failures of SVM in both Table I and Table II, 35% were perception errors and 65% were execution errors (see Figure 7). Within perception, mis-detections were surprisingly the dominant failure mode, despite using a vision foundation model. Within execution, insufficient precision was the dominant cause of failure. This suggests that future work should focus on strengthening eye-in-hand perception models and improving precise control under limited degrees of freedom.
@@ VI Discussion
In this paper, we describe SVM, a framework for precise manipulation tasks that involve reaching a precise interaction site followed by execution of a primitive motion. Vision foundation models help us mitigate issues caused by occlusion by the end-effector, thereby enabling the use of off-the-shelf open-vocabulary object detectors and point trackers to estimate servoing targets during execution. The use of strong off-the-shelf models also provides broad generalization for perception, while the use of servoing provides robust control. This enables SVM to solve tasks in novel environments on novel objects, obtaining a 71% zero-shot success rate across 6 precise mobile manipulation tasks. Perhaps most surprisingly, despite being modular, SVM outperforms the strong end-to-end imitation learning system RUM [7] which was trained on over 1000 demonstrations. This is particularly striking as imitation learning is often the tool of choice for precise tasks requiring closed-loop execution.
@@ VII Limitations
VIII Acknowledgements
This material is based upon work supported by DARPA (Machine Common Sense program), an NSF CAREER Award (IIS-2143873), and the Andrew T. Yang Research and Entrepreneurship Award. We are grateful to the Centre for Autonomy for lending us the Stretch RE2 robot used in this work. We thank Aditya Prakash and Xiaoyu Zhang for their feedback on manuscript.